# Claim Check-Worthiness Detection: How Well do LLMs Grasp Annotation Guidelines?

### v1

@@ 4.1 Prompt verbosity
Table 2: F1 scores of gpt-4-turbo for CD and CW tasks across datasets and prompt verbosity levels (V1–V3). Level V0 corresponds to the naive-prompting baseline.
4.2 Incorporating context
Claims are never made in isolation; their context matters not only for verifying their veracity but also for gauging their factuality and worthiness even before fact-checking. We investigate how LLMs’ predictive accuracy depends on the amount of situational context provided to the model. To this end, we leverage the context information available in the CB and POLI datasets and expand the prompts by adding the co-text of the claim (Level C1), speaker information (Level C2), or both (Level C3) (NEWS includes additional information, cf. Appendix A.3). The amount of co-text included in the prompt is the same as what was originally shown to the annotators (cf. Table 1). Speaker information pertains to the speaker’s identity and political party. The context information was appended to the prompts (cf. Appendix A for a detailed description).
We only extend the prompts with optimal verbosity level (V3 for CB and V1 for POLI). Table 3 shows the F1 scores by context levels for gpt-4-turbo. Surprisingly, expanding the prompt with co-text and speaker information did not improve the model’s accuracy on either dataset. This could be because context information is not required for these datasets, or because our specific prompt structure might be suboptimal and confuse the model. We leave further investigation for future work.
@@ 4.2 Incorporating context
C0 .919 .927 .844
C1 .877 .874 .790
C2 .906 .913 .707
C3 .846 .843 .692
Table 3: F1 scores of gpt-4-turbo by level context information added to the prompt (C1–C3). Level C0 corresponds to the prompt of optimal verbosity level with no context information (V3 for CB and V1 for POLI).
@@ 4.3 Rank-based evaluation
In light of resource constraints, fact-checking organizations have devised principles to prioritize claims based on their check-worthiness. This invites the question of whether zero- and few-shot LLM prompting could be used for that purpose. To investigate this, we frame CW as a ranking task and rank the claims based on the LLM’s confidence for the positive class. We used token likelihood of the positive class as a measure of confidence. The quality of the so-obtained ranking will depend on how well the LLM is calibrated. Thus, we first evaluate the LLMs’ calibration accuracy using the expected calibration error (ECE). Figure 1 shows the predictive accuracy (F1 score) against calibration accuracy ($1-\mathrm{ECE}$) across datasets and prompt verbosity levels (we only use prompts at context level C0, i.e., we add no context information). For each dataset, we select the prompt that scores high on both predictive and calibration accuracy. The prompts with the highest F1 scores are usually also the best-calibrated ones, except for NEWS, where we select level V1 as Pareto-optimal.
Figure 1: F1 scores and calibration accuracy for the CW task, across datasets and levels of prompt verbosity
Table 4 shows the rank-based performance scores for the selected prompts on gpt-4-turbo: average precision (AP), precision-at-10 (P@10), and precision-at-R, where R equals the total number of positives in the dataset. The rank-based performance scores mirror the classification accuracy scores: they are high for datasets with high predictive accuracy (CB and ENV) and lower for datasets with lower predictive accuracy (NEWS and CLEF). Our results suggest that LLM models with high predictive accuracy also produce well-calibrated scores using ECE and may be readily used as check-worthiness rankers.
@@
AP .951 .469 .767 .67
P@10 1 .5 .9 1
P@R .924 .487 .761 .533
Table 4: Rank-based performance scores for the CW task: average precision (AP), precision-at-10 (P@10), and precision-at-R (P@R)
@@ 5 Conclusion
We addressed the claim detection and check-worthiness tasks using zero- and few-shot LLM prompting based on existing annotation guidelines. The optimal level of prompt verbosity, ranging from minimal prompts to detailed prompts that include criteria and examples, depends on the domain and style of guidelines. Adding claim context (co-text and speaker information) does not improve the performance. For models with high predictive accuracy, confidence scores can be directly used to produce reliable check-worthiness rankings.
@@
In our experiments, we do not use datasets created by fact-checking organizations. While the datasets were created specifically for the tasks of CD and CW, and most were annotated by experts, the datasets were constructed for research purposes. To most accurately evaluate the potential of using our approach in fact-checking organizations, a dataset annotated according to official factuality or check-worthiness criteria with appropriate annotation guidelines should be used.
@@
Due to hardware constraints, no open-source LLMs of a larger size were used in our experiments. We acknowledge the importance of relying on open-source models in the research community, and the lack of insight that results from disregarding larger open-source models. Using closed-source models has the additional caveat of possible leakage of the dataset, which is a growing concern in the community (Balloccu et al., 2024). We also note that the outstanding results on the ClaimBuster dataset (CB) could be due to data leakage, considering the dataset was published several years ago and has a wide reach in the research of automatic fact-checking.
@@
CLEF&NEWS.
The results of weak performance on the CLEF and NEWS datasets could lie in the worthiness criteria used and the way the criteria are articulated. In CLEF, a Likert scale was used for the annotation. However, the levels in the scale do not completely correspond to gradation, as is usually the case. The negative labels include both tweets that do not need fact-checking (the label “No, no need to check”) and those worth fact-checking but not requiring experts’ attention (the label “No, too trivial to check”). This distinction probably creates ambiguity for the model, as demonstrated by poor predictive accuracy and calibration. On the NEWS dataset, the CW task essentially amounts to topic classification. It is unclear how the model should handle sentences unrelated to the provided topics. The authors in the corresponding research paper report performance scores with F1 not exceeding .7, even for annotators.
@@ Appendix A Dataset Information
In this section, we provide details on the datasets used in our experiments. In
@@ A.1 Test set selection
A.2 Original metrics
ClaimBuster.
As previously mentioned, the authors used 4-fold cross-validation on different-sized subsets during their experiments (4,000, 8,000 … 20,000). They evaluate using $f_{w}avg$ with the highest score of .818. Our highest scores on $f_{w}avg$ are .933 on gpt-4-turbo and .906 on gpt-3.5-turbo, which is a significant improvement. The authors also evaluate ranking, where our results improve on P@k.
CLEF.
The best results of the Task 1 on CLEF CheckThat!2022 were accuracy of .761 for claim detection, and the F1 of .698 on check-worthiness detection. While our approach underperforms for check-worthiness detection (F1 of .583), it achieves higher accuracy for claim detection (.776 on Level V2).
EnvironmentalClaims.
The authors report F1 on the test set, the highest achieved is .849. Our approach achieves .773 for Level V1.
NewsClaims.
The authors report F1 on the whole dataset. However, it is inconclusive whether it is evaluated on the binary or multiclass labels. They report the highest F1 of .309, which our approach surpasses on the subset we selected, with F1 of .583.
PoliClaim.
The authors evaluate on accuracy. On the test set, they achieve an accuracy of .764 on gpt-3.5 and .862 on gpt-4. Our approach achieves the same maximum accuracy using gpt-3.5-turbo with prompt Level V3, but lower accuracy for gpt-4.
A.3 Context information
@@
No additional contextual or co-textual information was provided in the dataset. The annotators were not shown any co-text during annotation due to budget. The authors considered annotating whole paragraphs instead of sentence-level annotation but decided against it due to time and budget constraints.
@@
The annotators were provided with the preceding and following sentences of the one they are annotating. Since there is only one speaker (as opposed to ClaimBuster, which covers debates), there is no need for denoting the speaker, minimizing confusion in prompts. In annotation guidelines, context was explicitly mentioned, as well as clarified in examples. In our experiments, we used two versions of the prompts – one mentioning context for experiments with co-text expansion and one without the mention of context used when only one sentence from the speech is provided. The two alternatives are shown in E.
@@
Appendix B Alternative Results
In Table 5 we present the results of prompt expansion for gpt-3.5-turbo-0125. In general, gpt-4-turbo outperforms this model, except for POLI, where Level V3 on gpt-3.5-turbo achieves the highest F1 score of .803.
@@ Appendix B Alternative Results
V0 .853 .718 .656 .496 .484 .531 .707
V1 .57 .739 .745 .438 .71 .371 .751
V2 .774 .800 .719 .468 .701 .348 .657
V3 .872 .862 .757 .445 .65 .206 .803
Table 5: F1 scores of gpt-3.5-turbo for CD and CW tasks across datasets broken down by prompt verbosity level (V1–V3). Level V0 corresponds to the naive-prompting baseline.
Appendix C Model Information
The initial experiments on open-source models were done on neural-chat:7b-v3.3-q5_K_M and mistral:7b-instruct-v0.2-q5_K_M. A total of 5 GPU hours was used.
Appendix D Calibration
In this section, the $ECE$ per prompt verbosity level is shown for gpt-3.5-turbo (Table 7), and gpt-4-turbo (Table 6). The $ECE$ is calculated with the parameters $n_{b}ins=10$ and $norm=l1$.
CB CLEF POLI ENV NEWS
CD CW CD CW CD CW CW
V0 .094 .068 .259 .601 .231 .322 .142
@@ Appendix D Calibration
Table 7: ECE score by prompt level per dataset for gpt-3.5-turbo. ’CD’ and ’CW’ mark claim detection and claim check-worthiness detection, respectively, while ’V0’ marks the score for the naive baseline
Appendix E Complete prompts

### v2

@@ 4.2 Amount of Context
Table 3: Binary F1 scores across datasets and prompt verbosity levels (V1–V3). Level V0 corresponds to the naive-prompting baseline. For baselines and previous results: a = accuracy, = not directly comparable
Level C1
represents adding the co-text of the claim. The amount of co-text included in the prompt for each dataset is the same as what was originally shown to the annotators – for CB, four preceding statements (which were either by the speaker, opposing speaker, or moderator), and for POLI, one preceding and one following statement;
Level C2
expands the contextual information by adding metadata to the claim. In the case of POLI, the metadata is the speaker’s identity and political party, whereas for CB it additionally contains the speaker’s title and the sentiment of the statement, provided by the authors of the dataset;
Level C3
combines both C1 and C2 by providing both co-text and metadata.
We appended the contextual information to the user prompts, and only modified the system prompts of POLI slightly – adding guidance on how to handle context, ommited from the no-context variants (cf. Appendix A for a detailed description).
5 Results
We present the results for prompt verbosity levels in Table 3 and for different context levels in Table 6. In Table 3, we also include the previous results reported by authors in the original papers introducing the datasets (note that some results are not directly comparable to ours, as we discuss below). We use a stratified random classifier, an SVM classifier with TF-IDF features, and a fine-tuned BERT (Devlin et al., 2019) as baselines.
5.1 Prompt Verbosity
Table 3 shows the baselines and F1 scores by verbosity level for gpt-4-turbo, gpt-3.5-turbo, and Llama3 8B. Both performance and the optimal verbosity level is not consistent across datasets. The accuracy generally increases with verbosity levels for CB, but the trend is reversed for ENV. We observe no consistent trend for CLEF, POLI, and NEWS datasets. The most verbose prompts (V3) generally do not achieve the highest performance, except for the GPT models and CB. This highlights that providing detailed instructions and examples can be beneficial but potentially harm performance.
Comparison to baselines.
For SVM and BERT baselines, we had to use a different test set for NewsClaims than for the other models, as the original dataset does not provide a training set (the best-performing LLM on this test set achieved an F1 score of 0.670). Overall, all best-performing LLMs outperform the SVM baseline. However, except for CLEF, BERT outperforms the best-performing LLMs by a small margin (<0.05 F1), raising doubts about whether manual data labeling is worthwhile in these cases.
Comparison to previous work.
Comparing with previous work is difficult due to differences in setup. For CB, the authors evaluated used 4-fold cross-validation on different-sized subsets (4,000, 8,000 …20,000), all containing our chosen test set, annotated by experts. The authors evaluated using weighted F1-score, achieving a maximum score of 0.818. Our highest weighted F1-scores surpass this, reaching 0.933 for gpt-4-turbo and 0.906 for gpt-3.5-turbo. On CLEF, the best-reported result is the accuracy score of .761 for CD and the F1 score of .698 for CW. While our approach underperforms for CW (F1 of 0.583), it achieves higher accuracy for CD (0.776 on Level V2). In the case of NEWS, the authors reported an F1 score, but it remains unclear whether it was evaluated based on binary or multiclass labels, given that annotators had to categorize claims into different classes. They achieved the highest F1 score of 0.309, which our approach exceeds on the subset we selected, achieving an F1 score of 0.583. Our subset has a higher random baseline due to a higher ratio of positive examples and includes all positives from the original test set. For POLI, the authors evaluated using accuracy. They achieved an accuracy of 0.764 on the test set using gpt-3.5 and 0.862 using gpt-4. The GPT-3.5-turbo using prompt Level V3 performs comparable, while GPT-4 and Llama3 perform worse. For ENV, the metrics are directly comparable, and our approach underperforms compared to previous results.
CB CLEF ENV NEWS POLI
CD CW CD CW CW CW CD
V1 26 9 8 55 13 150 87
V2 4 3 5 35 8 90 87
V3 1 1 2 14 5 68 65
Table 4: Counts of instances misclassified by all three models across the three verbosity levels.
Dataset Label Gold Example CB 1 0 Let’s go to work and end this fiasco in Central America, a failed policy which has actually increased Cuban and Soviet influence. CLEF 1 0 So businesses will get fined \$14,000 (per employee) if they don’t comply with Biden’s vaccine mandate. And illegal aliens get a \$450,000 payout for "damages" for crossing our border illegally… Biden’s America. ENV 1 0 Renewable energy is purely domestic sourced and environment-friendly, and can be used continuously without being depleted. 0 1 To strengthen its approach, Kering’s SBT for a 1.5°C trajectory was revised and approved by the SBTi in early 2021. NEWS 1 0 There’s currently no strong evidence that supplementing with vitamin C will prevent or cure COVID-19. 0 1 Vaccines, by their nature, are reactive. POLI 0 1 We are finally going to fix the darn roads. 0 1 Too many people are struggling to make ends meet.
Table 5: Examples of characteristic instances per dataset that were consistently misclassified across levels.
@@ 5.1 Prompt Verbosity
V1 V2 V3 V1 V2 V3 V1 V2 V3
gpt-4-turbo C0 .883 .908 .919 .885 .889 .927 .619 .541 .563
C1 .806 .849 .862 .803 .847 .872 .722 .650 .727
C2 .879 .908 .913 .880 .901 .916 .707 .470 .592
C3 .794 .857 .877 .791 .854 .885 .692 .632 .732
gpt-3.5-turbo C0 .570 .774 .872 .739 .800 .862 .751 .657 .803
C1 .461 .299 .513 .517 .301 .528 .790 .688 .794
C2 .560 .801 .836 .747 .826 .832 .730 .523 .704
C3 .474 .724 .758 .643 .716 .749 .794 .754 .800
Llama3 8B C0 .478 .742 .702 .655 .751 .637 .827 .712 .651
C1 .460 .591 .614 .531 .528 .552 .799 .789 .803
C2 .483 .773 .764 .727 .819 .736 .807 .703 .628
C3 .468 .610 .601 .506 .618 .556 .806 .798 .805
Table 6: Binary F1 scores by level of context information (C1–C3) added to the prompt ranging in verbosity (V1–V3) . Level C0 corresponds to the prompt level with no context information. The best scores across verbosity levels are shown in bold, and the best scores per model and dataset are highlighted in green.
CD vs. CW.
Generally, higher performance is achieved for the CD task, although the diverse domains of the datasets and differences in guidelines prevent definitive conclusions. Therefore, comparing performance on the two datasets that cover both tasks – CB and CLEF – is most straightforward. Interestingly, a reverse phenomenon is observed between these datasets—significantly higher performance is achieved for the CD task on CLEF, whereas on CB, CW performance is slightly higher. An important difference in the two datasets is precisely in the annotation styles – CB uses the same guidelines for both tasks and ternary annotation, while for CLEF the guidelines are different for the two tasks, originally using different labelling strategies (binary for CD and Likert scale for CW).
Closed-source vs. open-source.
While broader conclusions require a wider range of both open- and closed-source models, especially larger open-source ones, the Llama3 8B model performs similarly to GPT models, highlighting the potential of prompting open-source models with annotation guidelines. Furthermore, the results of both GPT models on the CB dataset could indicate a potential data leakage Balloccu et al. (2024) since the performance of Llama3 8B is comparable in other datasets but lags for CB.
5.2 Error Analysis
Worst performance.
The naive baseline prompt (V0) generally outperforms the prompts based on annotation guidelines on the CLEF CW and NEWS datasets, except for V2 for CLEF CW with gpt-4-turbo. For CLEF CW, the annotation guidelines are adapted from the Likert scale, where multiple characteristics are attributed to negatives (e.g., not interesting, a joke, not containing claims, or too trivial to be checked by a professional). In our prompts, we converted the Likert scale to binary, where the already diverse and vaguely defined criteria were binned in a single label, increasing complexity. For NEWS, although the dataset’s purpose is claim check-worthiness detection, check-worthiness as a concept is not mentioned in the annotation guidelines. Positives are merely selected by containing claims falling into four predefined categories relating to the COVID-19 virus, and check-worthiness is assumed implicitly. This, along with the presence of inter-sentence coreference in the positive instances, might cause poor performance.
Most difficult instances.
To analyze poor performance beyond the F1 score, we decided to identify the instances for which all three models across levels consistently predicted the wrong label. Table 4 shows the counts of those instances.
Interestingly, whether the instances in question were consistently misclassified as positives or negatives depends on the domain – for CLEF and CB, all of the instances are false positives (FP), whereas for POLI, all 65 are false negatives (FN). This suggests that the guidelines are too restrictive regarding the positive label or are interpreted as such during inference.
Table 5 shows examples from the final pool of mislabelled instances. Several interesting observations can be made here. For example, CLEF comprises tweets, where sarcasm is more prevalent, making the prediction task harder. For ENV, which contains environmental reports requiring expert knowledge, the mislabeled instances were either too vague for positives or contained too much domain knowledge for the models to decipher. Annotators were urged to look up acronyms they were not familiar with, but the same could not be accomplished with ICL. For NEWS, some claims are only implicitly related to COVID-19, which results in a false negative label. On the other hand, some instances seem mislabeled in the gold set. Concerning POLI, the frequent use of personal pronouns requiring coreference resolution leads to false negatives, with the claim losing relevance without information about the claimant.
5.3 Amount of Context
Table 6 shows the F1 scores by verbosity and context level for all models. The benefit of including context varies across models – there is a bigger performance increase for the Llama model with added contextual information, topping the performance for CB in both tasks as opposed to prompts with no context. For the GPT models, there is some positive impact of metadata (C2). The least beneficial is the addition of co-text but no metadata (C1), including speaker information, which is vital when given previous responses. Concerning prompt verbosity levels, context’s impact is higher on less verbose prompts, showing contextual information complements brief definitions.
@@ 5.4 Rank-Based Evaluation
In light of resource constraints, fact-checking organizations have devised principles to prioritize claims based on their check-worthiness. This invites the question of whether zero- and few-shot LLM prompting could be used for that purpose. To investigate this, we frame CW as a binary relevance ranking and rank the claims based on the LLM’s confidence for the positive class. We used the token likelihood of the positive class as a measure of confidence. The quality of the so-obtained ranking will depend on how well the LLM is calibrated. Thus, we first evaluate the LLMs’ calibration accuracy using the expected calibration error (ECE). Figure 2 shows the predictive accuracy (F1 score) against calibration accuracy ($1-\mathrm{ECE}$) across datasets and prompt verbosity levels (we only use prompts at context level C0, i.e., no context information).
@@
gpt-4 AP .951 .552 .767 .67
P10 1 .9 .9 1
PR .924 .615 .761 1
gpt-3.5 AP .934 .464 .796 .669
P10 1 .6 .9 .7
PR .919 .436 .772 .700
Llama3 8B AP .878 .350 .794 .688
P10 1 .2 .1 1
PR .823 .282 .762 1
Table 7: Rank-based CW performance scores
Figure 2: F1 scores and calibration accuracy ($1-\mathrm{ECE}$) for the CW task, across datasets and models
Per model and dataset, we select the prompt that scores high on predictive and calibration accuracy. The prompts with the highest F1 scores are usually the best-calibrated ones, except for NEWS, where we select level V1 as Pareto-optimal.
Table 7 shows the rank-based performance scores for the selected prompts: average precision (AP), precision-at-10 (P@10), and precision-at-R, where R equals the total number of positives in the dataset. The rank-based performance scores mirror the classification accuracy scores: they are high for datasets with high predictive accuracy (CB and ENV) and lower for datasets with lower predictive accuracy (NEWS and CLEF). Our results suggest that LLM models with high predictive accuracy also produce well-calibrated scores using ECE and may be readily used as check-worthiness rankers.
@@ 6 Conclusion
We tackled claim detection and check-worthiness tasks using zero- and few-shot LLM prompting based on existing annotation guidelines. The optimal level of prompt verbosity, from minimal prompts to detailed prompts that include criteria and examples, varies depending on the domain and guidelines style. Adding claim context does not improve performance. Models with high predictive accuracy can directly utilize confidence scores to produce reliable check-worthiness rankings.
@@
Our experiments do not use datasets created by fact-checking organizations. While the datasets were created specifically for the tasks of CD and CW, and most were annotated by experts, the datasets were constructed for research purposes. To most accurately evaluate the potential of using our approach in fact-checking organizations, a dataset annotated according to official factuality or check-worthiness criteria with appropriate annotation guidelines should be used.
@@
Due to hardware constraints, no open-source LLMs greater than 8B parameters were used in our experiments. We acknowledge the importance of relying on open-source models in the research community and the lack of insight that results from disregarding larger open-source models. Using closed-source models has the additional caveat of possible leakage of the dataset, which is a growing concern in the community (Balloccu et al., 2024). We also note that the outstanding results on the ClaimBuster dataset (CB) could be due to data leakage, considering the dataset was published several years ago and has a wide reach in the research of automatic fact-checking.
@@
Lack of prompt engineering experiments.
In this work, we do minimal prompt engineering interventions beyond merely adapting the level of detail in annotation guidelines and appending contextual information. We opted for this approach instead of drafting prompts ourselves to investigate how original wording, definitions, and examples given to annotators could fare with LLMs. We realize weak performance in some cases (e.g., CLEF, for the naive aggregation from the Likert scale to binary labels), and performance variations could be due to the models’ sensitivity to prompt structure, wording, and examples. However, translating the complex criteria of worthiness in such a streamlined way could benefit fact-checkers. Furthermore, prompt design should be adapted for each dataset, significantly expanding the scope of this research (since five datasets are used). We leave experiments regarding prompt design for future work.
Binary relevance ranking.
In our study, we also assess CW as a ranking task, which is crucial given that fact-checking organizations often face time and resource constraints and must prioritize claims based on their CW criteria. We use binary relevance judgments for evaluation, with binary CW labels as the ground truth, and apply standard IR metrics such as MAP and P@R. An alternative approach could involve using graded CW criteria, framing the task as regression, and evaluating with graded relevance judgments like NDCG. However, to our knowledge, only the dataset by Gencheva et al. (2017) provides graded CW labels, using aggregated judgments from various fact-checking agencies to model priority.
@@
Acknowledgments
This research was supported by the Adria Digital Media Observatory (ADMO) project, which is part of EDMO, EU’s largest interdisciplinary network for countering disinformation.
@@ Appendix A Dataset Information
In this section, we provide details on the datasets used in our experiments.
@@ A.1 Test set selection
A.2 Context information
@@
No additional contextual or co-textual information was provided in the dataset. The annotators were not shown any co-text during annotations. The authors considered annotating whole paragraphs instead of sentence-level annotation but decided against it due to time and budget constraints.
@@
The annotators were provided with the preceding and following sentences of the one they are annotating. Since there is only one speaker (as opposed to ClaimBuster, which covers debates), there is no need to differentiate the speaker to minimize confusion in prompts. In annotation guidelines, context was explicitly mentioned and clarified in examples. In our experiments, we used two versions of the prompts – one mentioning context for experiments with co-text expansion and one without the mention of context used when only one sentence from the speech is provided. The two alternatives are shown in D.
@@
Appendix B Model Information
We use Llama3 8B Instruct for experiments on open-source models. It is the only smaller open-source model from the ones we tested compliant with provided labels. The experiments took 10 GPU hours on 2x GeForce RTX 2080 Ti. We use greedy decoding and run once per model and prompt variant. Initial experiments were done on neural-chat:7b-v3.3-q5_K_M and mistral:7b-instruct-v0.2-q5_K_M. A total of 5 GPU hours was used.
For BERT, we use the base model bert-base-uncased. We train the model for five epochs with a batch size of 16, a learning rate of 2e-5 and weight decay of 0.01. We keep the best model across epochs.
Appendix C Calibration
In this section, the $ECE$ per prompt verbosity level is shown for all models in Table 8. The $ECE$ is calculated with the parameters $n_{b}ins=10$ and $norm=l1$.
@@ Appendix C Calibration
gpt-4-turbo V0 .094 .068 .259 .601 .231 .322 .142
@@
Llama3 8B V0 .218 .126 .307 .611 .286 .314 .223
V1 .607 .218 .244 .723 .114 .228 .172
V2 .184 .135 .241 .687 .102 .229 .321
V3 .231 .259 .241 .686 .134 .214 .379
Appendix D Complete prompts